Log[5](2x+10-log[5](x)=4
Answer by nerdybill(6962)

I think you forgot a right parenthesis. I'll assume you meant:
Log[5](2x+10)-log[5](x)=4
applying log rules:
Log[5](2x+10)/(x)=4
(2x+10)/(x)= 5^4
(2x+10)/x = 625
2x+10 = 625x
10 = 623x
10/623 = x
0.0161 = x